# On The Fragility of Benchmark Contamination Detection in Reasoning Models

### v1

@@ Appendix B Algorithms
  • Dataset Collection. For each question $q$, we sample a group of outputs $\{o_{1},\cdots,o_{G}\}$. For each response $o_{i}$, we compute the reward $r_{i}\in\{0,1\}$. We then retain only the responses that receive a reward of 11 and put them in a dataset $\mathcal{D}$.
@@ D.1 Contamination Pipelines
We randomly select 10K samples from OpenThoughts3 (Guha et al., 2025) to form the clean SFT training set, following the same ratio for each domain as the original paper, to obtain the best results. For the SFT contaminated training set, we use QwQ-32B (Team, 2025) as an advanced LRM to help distillation for the member set. We adopt the temperature 0.6, top-p value of 0.95, and maximum output token with 32768 for the distillation. We use rejection sampling with 64 rollouts, selecting correct trajectories when available. If none of the 64 rollouts produce a correct answer, we randomly select an incorrect trajectory. We replicate the question in the member set with responses 3 times to create the SFT contamination training set. So the training set consists of 11866 samples.
@@
We choose Qwen2.5-7B-Instruct (Team, 2024) as the base model, and first train it with SFT and then RL.
@@ Appendix E More Experiment Results
E.1 AUROC on Question, Question+Response, Thinking Process, and Non-thinking Process
We compare the AUROC computed on response tokens with computed on the question tokens, whole tokens, thinking tokens, and non-thinking tokens in the response. As shown in Tab. 7 and  10, none of the approach outperform the AUROC computed on the response tokens. Thus, we choose to compute all the detection scores on the response if applicable.
@@ E.1 AUROC on Question, Question+Response, Thinking Process, and Non-thinking Process
E.2 Log-probability distribution before and after RL
We provide the log probability distribution for members vs. non-members in diverse scenarios. Fig. 56, and 7 demonstrate that further SFT and RAFT are unable to contract the distribution for members and non-members, while GRPO and RAFT++ could conceal the contamination evidence due to the PPO-style importance sampling/clipping term in the training objective.
@@ E.2 Log-probability distribution before and after RL
E.3 AUROC for different training steps
We provide complete results of AUROC in different RL training steps. As shown in Tab. 11 and 12, we observe a monotonic decline in detection performance as the number of RL steps increases when using GRPO. While even with 156 steps in RAFT, there is no sign of AUROC decline. These results perfectly validate our theoretical analysis that RAFT is unable to conceal contamination, while GRPO could.
Table 11: AUROC (%) of contamination detection approaches evaluated starting from an SFT-contaminated model w/o RL to subsequently trained with GRPO in different steps. $\Delta$ measures the difference with the SFT contaminated model w/o RL (Tab. 2). The results demonstrate that additional GRPO steps better conceal the contamination evidence.
@@ E.3 AUROC for different training steps
Table 12: AUROC (%) of contamination detection approaches evaluated starting from an SFT-contaminated model w/o RL to subsequently trained with RAFT in different steps. $\Delta$ measures the difference with the SFT contaminated model w/o RL (Tab. 2). Results demonstrate that even with more RL steps, RAFT is unable to conceal the contamination evidence.
@@
E.4 Contamination Inflation Comparison

### v2

@@ Appendix B Algorithms
  • Dataset Collection. For each question $q$, we sample a group of outputs $\{o_{1},\cdots,o_{G}\}$. For each response $o_{i}$, we compute the reward $r_{i}\in\{0,1\}$. We then retain only the responses that receive a reward of $1$ and put them in a dataset $\mathcal{D}$.
@@ D.1 Contamination Pipelines
We randomly select 10K samples from OpenThoughts3 (Guha et al., 2025) to form the clean SFT training set, following the same ratio for each domain as the original paper, to obtain the best results. For the SFT contaminated training set, we use QwQ-32B (Team, 2025) as an advanced LRM to help distillation for the member set. We adopt the temperature 0.6, top-p value of 0.95, and maximum output token with 32768 for the distillation. We use rejection sampling with 64 rollouts, selecting correct trajectories when available. If none of the 64 rollouts produce a correct answer, we randomly select an incorrect trajectory. We replicate the question in the member set with responses 3 times to create the SFT contamination training set. So the training set consists of 11866 samples. For the Llama-3.1-8B-Instruct, we randomly select 30K samples from OpenThoughts3 (Guha et al., 2025) to form the clean SFT training set, and duplicate memberships 9 times to form the SFT contamination training set.
@@
We choose Qwen2.5-7B-Instruct (Team, 2024) and Llama-3.1-8B-Instruct Dubey et al. (2024) as the base model, and first train it with SFT and then RL.
@@ Appendix E More Experiment Results
E.1 More results of Llama-3.1-8B-Instruct (Stage I: pre-LRM)
Table 7: AUROC (%) of contamination detection approaches evaluated starting from an SFT-contaminated model w/o RL to subsequently trained with GRPO. Results demonstrate that after GRPO, AUROC decreases across all the benchmarks and detection approaches. $\Delta$ measures the difference with the SFT-contaminated model w/o RL. Higher AUROC, better detection performance. Each AUROC is averaged over detection scores from 8 rollouts. The base model is Llama3.1-8B-Instruct.
Generation based
Verbatim (Wu et al., 2025) Before RL 48.67 65.36 51.78 60.00 56.84 52.62 55.88 +0.00
RL w/ Clean 49.92 50.14 57.78 58.22 57.21 57.38 55.11 -0.77
RL w/ Clean&Mem 48.42 48.21 54.89 56.22 57.83 58.00 53.93 -1.95
CDD (Dong et al., 2024) Before RL 56.72 54.55 53.33 55.11 60.03 66.75 57.75 +0.00
RL w/ Clean 53.14 55.69 60.00 46.22 56.31 64.00 55.89 -1.86
RL w/ Clean&Mem 57.24 49.31 53.11 43.78 57.52 63.50 54.08 -3.67
Perturbation based
Neighbor (Mattern et al., 2023) Before RL 51.64 40.46 62.44 41.11 53.86 63.62 52.19 +0.00
RL w/ Clean 52.48 37.07 55.78 39.33 50.32 65.00 50.00 -2.19
RL w/ Clean&Mem 52.34 39.76 49.78 37.11 54.44 70.25 50.61 -1.58
Reference based
LiRA (Mireshghallah et al., 2022) Before RL 80.50 70.86 98.22 93.78 81.54 98.87 87.30 +0.00
RL w/ Clean 68.95 65.57 51.33 41.11 73.99 80.75 63.62 -23.68
RL w/ Clean&Mem 68.72 65.74 90.44 89.11 72.32 94.12 80.08 -7.22
Ref (Carlini et al., 2021) Before RL 60.21 48.62 66.00 40.89 60.47 83.50 59.95 +0.00
RL w/ Clean 55.27 47.41 40.67 32.44 52.31 53.25 46.89 -13.06
RL w/ Clean&Mem 55.49 50.37 46.00 31.11 52.98 62.50 49.74 -10.21
Reference free
Zlib (Carlini et al., 2021) Before RL 49.46 56.21 88.44 72.00 47.04 60.25 62.23 +0.00
RL w/ Clean 46.53 56.37 66.67 50.00 45.91 41.38 51.14 -11.09
RL w/ Clean&Mem 47.29 52.29 66.22 60.89 45.63 46.00 53.05 -9.18
Min–K%++ (Zhang et al., 2024) Before RL 48.88 49.33 41.43 39.11 51.67 30.63 43.51 +0.00
RL w/ Clean 46.67 47.61 29.33 24.22 59.85 34.38 40.34 -3.17
RL w/ Clean&Mem 48.08 50.65 36.22 28.44 59.27 46.63 44.88 +1.37
Min–K% (Shi et al., 2023) Before RL 68.90 72.17 98.22 99.56 76.16 94.50 84.92 +0.00
RL w/ Clean 59.11 69.85 60.44 67.56 73.89 75.87 67.79 -17.13
RL w/ Clean&Mem 60.41 63.87 87.33 89.56 73.57 84.50 76.54 -8.38
Max–K% (Maini et al., 2024) Before RL 63.60 61.58 92.44 89.33 69.12 91.50 77.93 +0.00
RL w/ Clean 50.59 52.50 50.00 50.00 50.37 52.50 50.99 -26.94
RL w/ Clean&Mem 50.91 52.24 49.11 49.11 54.86 61.75 53.00 -24.93
Loss (Carlini et al., 2021) Before RL 69.44 72.88 98.67 99.56 76.91 94.50 85.33 +0.00
RL w/ Clean 59.18 69.84 60.00 68.00 74.02 76.75 67.97 -17.36
RL w/ Clean&Mem 60.37 63.92 89.33 90.44 74.18 84.62 77.14 -8.19
[…]
E.2 AUROC on Question, Question+Response, Thinking Process, and Non-thinking Process (Stage I: pre-LRM)
We compare the AUROC computed on response tokens with computed on the question tokens, whole tokens, thinking tokens, and non-thinking tokens in the response. As shown in Tab. 8 and  11, none of the approach outperform the AUROC computed on the response tokens. Thus, we choose to compute all the detection scores on the response if applicable. The base model here is Qwen2.5-7B-Instruct.
@@ E.2 AUROC on Question, Question+Response, Thinking Process, and Non-thinking Process (Stage I: pre-LRM)
E.3 Response Length Analysis (Stage I: pre-LRM)
We visualize the average response length per question across different benchmarks in Fig. 6. Here, the base model is Qwen2.5-7B-Instruct. Although the average response length does not change too much after GRPO training, several detection approaches already exhibit a substantial performance drop, as shown in Tab. 2. In contrast, for RAFT, even though the response lengths remain similar to those after GRPO, no concealment effect is observed. These results suggest that response length is not the key factor; instead, the entropy of the model’s output appears to drive the concealment phenomenon.
Figure 6: Average response length for each question across different benchmarks.
E.4 Log-probability distribution before and after RL (Stage I: pre-LRM)
We provide the log probability distribution for members vs. non-members in diverse scenarios. Fig. 78, and 9 demonstrate that further SFT and RAFT are unable to contract the distribution for members and non-members, while GRPO and RAFT++ could conceal the contamination evidence due to the PPO-style importance sampling/clipping term in the training objective. The base model here is Qwen2.5-7B-Instruct.
@@ E.4 Log-probability distribution before and after RL (Stage I: pre-LRM)
E.5 AUROC for different training steps (Stage I: pre-LRM)
We provide complete results of AUROC in different RL training steps. As shown in Tab. 12 and 13, we observe a monotonic decline in detection performance as the number of RL steps increases when using GRPO. While even with 156 steps in RAFT, there is no sign of AUROC decline. These results perfectly validate our theoretical analysis that RAFT is unable to conceal contamination, while GRPO could.
Table 12: AUROC (%) of contamination detection approaches evaluated starting from an SFT-contaminated model w/o RL to subsequently trained with GRPO in different steps. $\Delta$ measures the difference with the SFT contaminated model w/o RL (Tab. 2). The results demonstrate that additional GRPO steps better conceal the contamination evidence. The base model here is Qwen2.5-7B-Instruct.
@@ E.5 AUROC for different training steps (Stage I: pre-LRM)
Table 13: AUROC (%) of contamination detection approaches evaluated starting from an SFT-contaminated model w/o RL to subsequently trained with RAFT in different steps. $\Delta$ measures the difference with the SFT contaminated model w/o RL (Tab. 2). Results demonstrate that even with more RL steps, RAFT is unable to conceal the contamination evidence. The base model here is Qwen2.5-7B-Instruct.
@@
Table 14: AUROC (%) of contamination detection approaches evaluated starting from an SFT-contaminated model w/o RL to subsequently trained with GRPO in different steps. $\Delta$ measures the difference with the SFT contaminated model w/o RL (Tab. 7). The results demonstrate that additional GRPO steps better conceal the contamination evidence. The base model here is Llama-3.1-8B-Instruct.
Generation base
Verbatim (Wu et al., 2025) 64 49.92 50.14 57.78 58.22 57.21 57.38 55.11 -0.77
110 49.53 51.25 54.22 60.67 54.73 53.12 53.92 -1.96
156 49.45 52.11 51.11 60.44 57.02 47.12 52.88 -3.00
CDD (Dong et al., 2024) 64 53.14 55.69 60.00 46.22 56.31 64.00 55.89 -1.89
110 53.48 48.40 46.67 49.11 57.34 56.00 51.83 -5.92
156 53.99 52.18 45.56 52.89 60.62 57.25 53.75 -4.00
Perturbation base
Neighbor (Mattern et al., 2023) 64 52.48 37.07 55.78 39.33 50.32 65.00 50.00 -2.19
110 52.83 38.31 49.78 39.56 50.18 60.00 48.44 -3.75
156 53.09 39.15 48.00 44.00 52.29 62.25 49.80 -2.39
Reference base
LiRA (Mireshghallah et al., 2022) 64 68.95 65.57 51.33 41.11 73.99 80.75 63.62 -23.68
110 66.57 65.72 44.22 27.56 72.02 71.00 57.85 -29.45
156 60.93 64.21 40.22 23.11 69.25 70.12 54.64 -32.66
Ref (Carlini et al., 2021) 64 55.27 47.41 40.67 32.44 52.31 53.25 46.89 -13.06
110 54.98 49.70 42.89 26.67 53.16 58.00 47.57 -12.38
156 53.23 47.75 43.11 27.33 56.31 56.75 47.41 -12.54
Reference free
Zlib (Carlini et al., 2021) 64 46.53 56.37 66.67 50.00 45.91 41.38 51.14 -11.09
110 46.78 57.11 64.89 48.44 46.40 37.25 50.15 -12.08
156 46.32 54.41 65.78 39.78 42.50 42.88 48.61 -13.62
Min–K%++ (Zhang et al., 2024) 64 46.67 47.61 29.33 24.22 59.85 34.38 40.34 -3.17
110 49.92 50.25 42.89 27.78 57.32 42.75 45.15 +1.64
156 49.42 48.28 41.33 38.67 53.09 48.25 46.51 +3.00
Min–K% (Shi et al., 2023) 64 59.11 69.85 60.44 67.56 73.89 75.87 67.79 -17.13
110 56.75 66.89 54.89 64.22 71.61 62.25 62.77 -22.15
156 55.83 65.12 50.44 59.78 62.76 62.50 59.41 -25.51
Max–K% (Maini et al., 2024) 64 50.59 52.50 50.00 50.00 50.37 52.50 50.99 -26.94
110 50.29 54.44 50.00 50.00 50.00 52.50 51.21 -26.72
156 50.30 51.41 53.33 50.00 51.09 50.00 51.02 -26.91
Loss (Carlini et al., 2021) 64 59.18 69.84 60.00 68.00 74.02 76.75 67.97 -17.36
110 56.85 66.90 55.11 63.78 71.60 62.38 62.77 -22.56
156 55.75 65.06 50.00 59.78 62.72 63.00 59.39 -25.94
Table 15: AUROC (%) of contamination detection approaches evaluated starting from an SFT-contaminated model w/o RL to subsequently trained with RAFT in different steps. $\Delta$ measures the difference with the SFT contaminated model w/o RL (Tab. 7). Results demonstrate that even with more RL steps, RAFT is unable to conceal the contamination evidence. The base model here is Llama-3.1-8B-Instruct.
Generation base
Verbatim (Wu et al., 2025) 64 46.75 49.07 56.44 58.89 55.77 57.00 53.99 -1.89
110 46.89 50.85 52.00 53.33 58.78 55.12 52.83 -3.05
156 47.06 49.70 54.22 60.44 55.63 61.75 54.80 -1.08
CDD (Dong et al., 2024) 64 57.35 54.44 52.22 46.44 53.57 62.62 54.44 -3.31
110 56.97 54.92 60.00 50.00 55.76 65.38 57.17 -0.58
156 55.78 57.84 53.11 50.89 60.63 70.88 58.19 +0.44
Perturbation base
Neighbor (Mattern et al., 2023) 64 51.81 40.93 64.00 48.44 55.28 66.50 54.49 +2.30
110 52.08 39.36 64.22 54.00 54.84 68.88 55.56 +3.37
156 51.99 40.25 66.22 49.56 51.68 72.75 55.41 +3.22
Reference base
LiRA (Mireshghallah et al., 2022) 64 81.82 77.88 96.89 96.67 82.92 97.50 88.95 +1.75
110 81.54 77.08 99.56 95.78 86.03 99.00 89.83 +2.63
156 80.52 77.21 99.11 96.00 84.04 97.75 89.11 +1.91
Ref (Carlini et al., 2021) 64 60.81 51.55 65.33 49.11 61.41 88.00 62.70 +2.75
110 62.25 49.94 76.89 52.67 61.33 86.00 64.85 +4.93
156 61.14 48.53 77.78 49.78 63.06 87.50 64.63 +4.68
Reference free
Zlib (Carlini et al., 2021) 64 53.02 58.98 80.67 68.44 51.21 60.75 62.18 -0.05
110 53.40 64.35 89.33 66.22 53.83 67.88 65.84 +3.61
156 54.13 62.78 85.33 64.89 54.81 68.62 65.09 +2.86
Min–K%++ (Zhang et al., 2024) 64 48.97 48.15 36.00 36.89 55.41 37.25 43.78 +0.27
110 51.00 56.82 44.44 38.89 58.25 39.47 48.15 +4.64
156 50.43 58.89 34.44 29.11 58.98 45.50 46.23 +2.72
Min–K% (Shi et al., 2023) 64 73.95 76.84 98.00 98.89 82.44 97.25 87.90 +2.98
110 71.87 80.78 98.89 98.22 83.92 95.88 88.26 +3.34
156 71.30 81.59 98.22 100.00 82.54 95.50 88.19 +3.27
Max–K% (Maini et al., 2024) 64 66.57 70.43 92.44 89.11 77.31 95.25 81.85 +3.92
110 67.37 75.39 92.89 92.00 78.65 96.88 83.86 +5.93
156 67.61 73.85 90.67 94.22 76.10 95.50 82.99 +5.06
Loss (Carlini et al., 2021) 64 73.77 78.18 98.00 99.56 83.14 97.38 88.34 +3.01
110 72.40 81.76 99.11 98.67 84.18 97.75 88.98 +3.65
156 71.57 82.91 98.67 100.00 82.34 95.75 88.54 +3.21
E.6 Contraction effect on small and large models (Stage I: pre-LRM)
We conduct experiments on smaller and larger models to see whether the concealment of RL has correlations with the model size. In particular, we provide additional results with a small model (i.e., Qwen2.5-3B-Instruct) in Table 16, and a large model (i.e., Qwen2.5-14B-Instruct) in Table 17. We compare them with the contraction of the medium-sized model (i.e., Qwen2.5-7B-Instruct). Here, RL steps are 64 by default. The empirical results further validate that the contraction extends to a large range of model sizes. Regardless of different model sizes, the RL training could still conceal the contamination introduced in the SFT stage.
Table 16: AUROC (%) of contamination detection approaches evaluated starting from an SFT-contaminated model w/o RL to subsequently trained with GRPO. Results demonstrate that after GRPO, AUROC decreases across all the benchmarks and detection approaches. $\Delta$ measures the difference with the SFT-contaminated model w/o RL. The base model here is Qwen2.5-3B-Instruct.
Reference based
LiRA (Mireshghallah et al., 2022) Before RL 54.89 48.60 88.67 95.56 61.74 90.25 73.29 +0.00
RL w/ Clean 53.93 51.17 62.89 81.56 60.10 81.50 65.19 -8.09
Ref (Carlini et al., 2021) Before RL 52.66 43.09 84.00 73.33 51.75 75.75 63.43 +0.00
RL w/ Clean 53.65 44.54 66.44 63.33 52.24 73.25 58.91 -4.52
Reference free
Loss (Carlini et al., 2021) Before RL 51.75 57.64 58.00 76.00 56.84 64.75 60.83 +0.00
RL w/ Clean 49.20 57.54 42.22 55.11 54.83 54.75 52.28 -8.56
Min–K (Shi et al., 2023) Before RL 50.86 57.24 51.11 73.33 56.47 59.13 58.02 +0.00
RL w/ Clean 48.71 56.93 39.56 49.78 54.90 51.37 50.21 -7.82
Max–K (Maini et al., 2024) Before RL 52.86 54.58 88.44 88.44 51.35 79.12 69.13 +0.00
RL w/ Clean 52.07 54.54 56.67 57.78 52.57 74.00 57.94 -11.19
Table 17: AUROC (%) of contamination detection approaches evaluated starting from an SFT-contaminated model w/o RL to subsequently trained with GRPO. Results demonstrate that after GRPO, AUROC decreases across all the benchmarks and detection approaches. $\Delta$ measures the difference with the SFT-contaminated model w/o RL. The base model here is Qwen2.5-14B-Instruct.
Reference based
LiRA (Mireshghallah et al., 2022) Before RL 87.35 86.54 91.56 85.33 88.12 91.25 88.36 +0.00
RL w/ Clean 83.46 82.38 90.56 84.00 82.23 84.88 84.59 -3.77
Ref (Carlini et al., 2021) Before RL 62.33 54.69 60.44 37.78 59.26 65.25 56.63 +0.00
RL w/ Clean 56.28 58.85 50.67 36.00 53.32 63.88 53.17 -3.46
Reference free
Loss (Carlini et al., 2021) Before RL 73.59 76.98 79.11 88.00 77.45 84.88 80.00 +0.00
RL w/ Clean 63.50 61.21 67.33 89.11 71.73 68.75 70.27 -9.73
Min–K (Shi et al., 2023) Before RL 73.04 76.16 77.11 86.22 77.61 83.75 78.98 +0.00
RL w/ Clean 63.15 60.09 66.67 87.78 71.69 67.12 69.42 -9.57
Max–K (Maini et al., 2024) Before RL 63.75 71.46 64.00 84.89 57.99 78.00 70.02 +0.00
RL w/ Clean 52.94 59.10 46.67 53.33 48.89 55.00 52.66 -17.36
E.7 More RL training brings performance gain (Stage I: pre-LRM)
In Table 1, we note that sometimes RL does not bring any performance gain, even in the clean setting. To clarify that our RL setting and recipe are correct, we would like to point out that the insignificant RL improvement was due to our relatively short RL training run to save experiment cost, which already effectively demonstrated the contamination concealment effect, whereas RL fine-tuning on reasoning models typically requires many more training steps to boost performance significantly (i.e, thousands of RL steps, more than 16k GPU hours, etc).
However, we further verified our RL pipeline by continuing RL training on the SFT-contaminated Qwen-3B model for up to 280 steps. The RL concealment still exists: the average AUROC of LiRA on the AIME24, AIME25, and AMC23 datasets drops from 91.49% before RL to approaching random guessing after 280 steps. At the same time, RL continues to improve task performance, as reported in Table 18. Overall, these results demonstrate that additional RL training on an SFT-contaminated model can both inflate performance and further conceal contamination signals.
Table 18: Pass@1 (%) across six reasoning benchmarks. Results show that additional RL training on an SFT-contaminated model can bring performance gain.
Model Olympiad GPQA AIME25 AIME24 Minerva AMC23 Avg.
Qwen2.5-3B-Instruct 26.98 31.06 0.83 6.67 21.69 36.07 20.54
   $\hookrightarrow$ SFT contamination 35.07 37.11 13.23 14.32 29.13 44.06 28.82
   $\hookrightarrow$ Further RL 40.74 38.38 15.12 15.83 31.25 49.06 31.73
E.8 Token Embedding Visualization of member/non-member of benchmarks (Stage I: pre-LRM)
We provide UMAP visualization of token embeddings of member and non member of benchmarks. Specifically, we input the entire question and response to the model, and extract the last token from the last layer of the hidden space. The base model is selected as Qwen2.5-7B-Instruct. We choose n_neighbors as 15, min_dist as 0.1, n_components as 2, and metric as cosine. As shown in Fig. 10,  11, and  12, the visualization indicates that member/non-member embeddings are highly overlapped and are hard to distinguish. We also provide quantitative evaluations of embedding based detection following (Liu et al., 2024) in Table 19. In particular, we randomly split all member and non-member examples into training and test sets, maintaining a consistent 0.8/0.2 ratio across datasets, and train an MLP classifier to predict the member/non-member labels. The input to the MLP is the hidden state feature of the last token in the question and response pair. Following (Liu et al., 2024), we evaluated several hidden layers and selected the best performing one. We report results only on the medium-scale benchmarks (Olympiad, GPQA, Minerva), since the small-scale datasets have very few test examples, resulting in too much variance. The results show that contamination is initially detectable, but becomes undetectable after RL.
Table 19: AUROC (%) of the embedding-based detection (Liu et al., 2024) evaluated starting from an SFT-contaminated model w/o RL to subsequently trained with GRPO. Results demonstrate that after GRPO, AUROC decreases. $\Delta$ measures the difference with the SFT-contaminated model w/o RL. Here, the base model is Qwen2.5-7B-Instruct.
Training Stage Olympiad GPQA Minerva Avg. $\Delta$
Before RL 58.89 80.00 60.05 66.31 +0.00
RL w/ Clean 56.63 45.75 49.60 50.66 -15.65
RL w/ Clean&Mem 57.43 45.00 48.94 50.46 -15.85
(a) Base model
(b) SFT contaminated model
Figure 10: Last token embedding visualization on the Minerva Math dataset.
(a) Base model
(b) SFT contaminated model
Figure 11: Last token embedding visualization on the GPQA-Diamond dataset.
(a) Base model
(b) SFT contaminated model
Figure 12: Last token embedding visualization on the OlympiadBench dataset
E.9 Contamination Inflation Comparison (Stage I: pre-LRM)